Calculate the approximate mass, in kg, of a 2.0m length of cylindrical clay pipe of external and internal diameters 14cm and 12cm. Density of clay is 1.3g/cm3.
Answer by mananth(16949)

Calculate the approximate mass, in kg, of a 2.0m length of cylindrical clay pipe of external and internal diameters 14cm and 12cm. Density of clay is 1.3g/cm3.
...
Outer Volume V1= pi*7^2*200= 30,787.60
Inner Volume V2= pi*6^2*200= 22,619.46
..
V1-V2= Volume of clay = 8168.13 cm^3
Density = 1.3 g/cm^3
Mass = 1.3 *8168.13 grams
Mass = 10,618,50 grams
10.6kgs
...
m.ananth@hotmail.ca